A recent article suggested that if you ear BDT 260000 a year today and the inflation rate continues at 3 percent per year, you’ll need to make BDT 335980 in 10 years to have the same buying power. You would need to make BDT 447710 if the inflation rate jumped to 6 percent. Please confirm whether these statements are accurate or not. Please show all related procedure behind your decision making process. 



1
Expert's answer
2020-12-08T11:18:13-0500

case 1

335980=260000*(1+x)10

1+x="^{10}\sqrt{335980/260000}"

1+x=1.026

x=0.026=2.6%"\neq3" %

first statement are not accurate

case 2

447710 =260000*(1+x)10

1+x="^{10}\sqrt{447710 /260000}"

1+x=1.056

x=0.056=5.6%"\neq6%" %

second statement are not accurate